Mon Jan 7  Tue Jan 8  Wed Jan 9 Thu Jan 10 Fri Jan 11  Sat Jan 12

07:00

08:30  Breakfast  Breakfast  Breakfast  Breakfast  Breakfast
08:30  W.Busza  G.Altarelli-1 J.Harris  F.Karsch  J. Harris
09:20  G. Altarelli-1  W.Busza  F.Karsch  F.Antinori  A.DiGiacomo
10:10  Break  Break  Break  Break  Break
10:40  T.Kodama  F.Karsch  W.Busza A.DiGiacomo  F.Antinori
11:30  ” ”  J.Harris  G.Altarelli  K.Safarik  K.Safarik

12:20

12:30  Lunch  Lunch  Lunch  Lunch  Lunch
14:00  CHECKIN  Discussion  Discussion Discussion  Discussion  Discussion
15:30  Self  J.Aichelin  J.Aichelin HADRONS U.VanKolck  U.VanKolck
16:20 Introduction  C.Bertulani  T.Trainor  IN LATIN  Ch.Roland  A.Szanto
17:00  Snack  Break  Break  –  Break  Break
17:30  Self  D. Kharzeev-1 B.Müller-1 AMERICA  L.Masperi  N. Scoccola
18:10 Introduction  ” ”  ” ”  ””  C.Vasconcelos  Student-1
18:35  ” ”  ” ”  ” ”  ””  M.Malheiro  Student-2
19:00  Reception
20:00  Dinner  Dinner  Dinner  Dinner  Dinner  Dinner
Sun Jan 13  Mon Jan 14 Tue Jan 15 Wed Jan 16 Thu Jan 17  Fri Jan 18

07:00

08:30 DEPART FOR  Breakfast  Breakfast  Breakfast  Breakfast  Breakfast
08:30 EXCURSION L.McLerran L.McLerran L.McLerran  B.Müller  B.Müller
09:20  G.Altarelli  G.Altarelli  H.T.Elze  H.T.Elze  K.Werner
10:10  Break  Break  Break  Break  Break
10:40  L.Csernai  L.Csernai  L.Csernai  D.Kharzeev  D.Kharzeev
11:30  H.Stöcker  H.Stöcker  K.Werner  R.Thews  R.Thews

12:20

12:30  BOX  Lunch  Lunch  Lunch  Lunch  Closing/Lunch
14:00  Lunch  Discussion  Discussion Discussion  Discussion  Dep:*14:30*
15:30  J.Rafelski  J.Rafelski  J.Rafelski  Student
16:20  H.Takai  G.Herrera  L.Turko  Program
17:00  Break  Break  Break  Break
17:30  A.Ayala  F.Grassi  M.Nielsen  C.Tsallis
18:10  M.G.Ducati  Student 4  T.Osada  ” ”
18:45  Student 3  Student 5  Student 6  ” ”
19:00  BBQ Dinner
20:00  Dinner  Dinner  Dinner  Dinner  ” ”

2

PASI-NEW STATES OF MATTER January 7–18, 2002

Lecture Courses: N × 40 +10 min.

1. W.Busza: 3 ×“Introduction to Relativistic Heavy Ion Collisions”

2. G.Altarelli-1: 3 ×“Introduction to Quantum Chromodynamics”

3. J.Aichelin: 2 × “Introduction to Hot Hadronic Matter”

4. F. Karsch: 3 × “Introduction to Lattice QCD”

5. J.Harris: 3 × “RHIC Physics - Overview”

6. A. DiGiacomo: 2 × “Vacuum Properties of QCD”

7. F. Antinori: 2 × “SPS Physics - Overview”

8. K.Safarik: 2 × “Introduction to Particle Tracking”

9. U.VanKolck: 2 × “Introduction to Chiral Symmetries and Models”

10. L.McLerran: 3 × “Color Glass Model and Structure Functions”

11. G.Altarelli-2: 2 ×“QCD Structure Functions”

12. L.Csernai: 3 × “Relativistic Hydrodynamics”

13. H-T. Elze: 2 × “Relativistic Quantum Transport Theory”

14. J.Rafelski: 3 × “Strangeness Signature of QGP”

15. H.Stöcker: 2 × “Critical Review of QGP Signals”

16. B.Müller-2: 2 × “Thermal QCD”

17. K.Werner: 2 × “Physics of Event Generators”

18. D.Kharzeev-2: 2 × “Charm Suppression”

19. R.Thews: 2 × “Charm Enhancement”

Tutorial Lectures: One double course lecture

1. D.Kharzeev-1: “Hadron-hadron and Hadron-nucleus Interactions”

2. T.Kodama: “Physics of Relativistic Gases”

3. B.Müller-1: “Quark Matter”

DISCUSSION: Recent course speakers to be available for discussions with iterested students, group discussios of speakers presentation write up.

Topical Lectures: 30+10 min.

1. C.Bertulani: “Peripheral Collisions”

2. T. Trainor: “Purpose and Physics of Event by Event Analysis”

3. Ch.Roland: “Survey of Recent PHOBOS Results”

4. L. Masperi: “Strangelets in the Universe”

5. A.Szanto: “Survey of Recent RHIC Results”

6. N. Scoccola: “Finite Temperature Chiral Models”

7. H.Taki: “More Heavy Ions at LHC”

8. A.Ayala: “Hyperon Polarization”

9. G.Herrera: “Charge Asymmetries in Charm Production”

10. L.Turko: “Non-Abelian Internal Symmetries”

11. F.Grassi: “Continuous Emission”

12. M.Nielsen: “QCD Sum Rules” Coloquium:

1. C. Tsallis: “Non-Extensive Statistics and Its Applications”

Mini-Symposium:

• “History and Development of Strong Interaction Physics in Latin America”, Coordinator: E.Ferreira Seminars: 20+5 min.

1. C.Vasconcelos: “Quark Stars - Revisited”

2. M.Malheiro: “Quark Matter in Chromodielectric Model”

3. M.G.Ducati: “Unitarity Corrections to Structure Functions”

4. T.Osada: “Event by Event Hydrodynamics”

5. Student 1

6. Student 2 7. Student 3 8. Student 4 9. Student 5 10. Student 6